OpenStudy (anonymous):

dibbie anderson a piece of wood measuring 2 1/8 inches into two equal pieces. howlong is each piece? the length of each piece is

OpenStudy (mertsj):

\[2 \frac{1}{8}\div 2=\frac{17}{8}\times \frac{1}{2}\]

OpenStudy (mertsj):

You can finish it.

OpenStudy (anonymous):

im lost

OpenStudy (mertsj):

If you divide a dollar into two equal parts, how big is each part?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

or 50 and 50

OpenStudy (mertsj):

If you have 1 dollar and you keep half of it and give half of it to your friend, how much do you each get?

OpenStudy (anonymous):

50

OpenStudy (mertsj):

And what is 100 divided by 2?

OpenStudy (anonymous):

200

OpenStudy (mertsj):

100 times 2 is 200 What is 100 divided by 2?

OpenStudy (anonymous):

50

OpenStudy (mertsj):

So the concept here is that if you want to divide something into two equal parts, you should divide by 2.

OpenStudy (mertsj):

For example if you have 12 eggs and want to put the same number of eggs into two containers, you should say: 12 divided by 2 is 6 so I should put 6 eggs in each container.

OpenStudy (mertsj):

It you have 10 tootsie rolls and want to give an equal amount to each of your two children you should say: 10 divided by 2 is 5 and you should give each child 5 tootsie rolls.

OpenStudy (mertsj):

If you have a piece of wood that is 2 1/8 inches long and you want to cut it so that you end up with two equal pieces you should say 2 1/8 divided by 2 is ??? and you should do that problem and find out how long each piece would be.

OpenStudy (primeralph):

dibbie anderson is a piece of wood?

OpenStudy (mertsj):

@primeralph Don't nit pick. She probably meant "has" instead of "a". Maybe English is a second language or something. Maybe it was a typo.

OpenStudy (primeralph):

@Mertsj To be honest, I didn't mean to. I was just asking. Farmers or such, sometimes give names to their tools and the likes.